# Implicit Actor Critic Coupling via a Supervised Learning Framework for RLVR

### v1

@@ Abstract
Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) have empowered large language models (LLMs) to tackle challenging reasoning tasks such as mathematics and programming. RLVR leverages verifiable outcome rewards to guide policy optimization, enabling LLMs to progressively improve output quality in a grounded and reliable manner. Despite its promise, the RLVR paradigm poses significant challenges, as existing methods often suffer from sparse reward signals and unstable policy gradient updates, particularly in RL-based approaches. To address the challenges, we propose PACS, a novel RLVR framework that achieves imPlicit Actor Critic coupling via a Supervised learning framework. By treating the outcome reward as a predictable label, we reformulate the RLVR problem into a supervised learning task over a score function parameterized by the policy model and optimized using cross-entropy loss. A detailed gradient analysis shows that this supervised formulation inherently recovers the classical policy gradient update while implicitly coupling actor and critic roles, yielding more stable and efficient training. Benchmarking on challenging mathematical reasoning tasks, PACS outperforms strong RLVR baselines, such as PPO and GRPO, achieving superior reasoning performance. For instance, PACS achieves 59.78% at pass@256 on AIME 2025, representing improvements of 13.32 and 14.36 points over PPO and GRPO. This simple yet powerful framework offers a promising avenue for LLMs post-training with verifiable rewards. Our code and data are available as open source at https://github.com/ritzz-ai/PACS.
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved remarkable success in complex reasoning tasks, driven by advancements, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). In particular, scaling test-time compute enables longer Chains of Thought (CoT), facilitating more sophisticated and structured reasoning. This progress has led to significant improvements in challenging domains, such as mathematics (Shao et al., 2024; Liu et al., 2025a; Zuo et al., 2025) and programming (Lambert et al., 2024; Cui et al., 2025; Zhao et al., 2025). A crucial underlying ingredient behind these achievements is the emergence of Reinforcement Learning with Verifiable Rewards (RLVR), which empowers LLMs to progressively enhance their capabilities by reinforcing their self-generated outputs with verifiable outcome rewards.
Existing RLVR methods fall into two primary categories: value-model-based approaches, such as PPO (Schulman et al., 2017) and VAPO (Yue et al., 2025), which employ large value models to estimate advantages for policy updates; and value-model-free approaches, such as GRPO (Shao et al., 2024; Guo et al., 2025) and DAPO (Yu et al., 2025), which bypass explicit value modeling by leveraging Monte Carlo estimation of group relative rewards. While both classes have demonstrated remarkable success, they face fundamental challenges stemming from the sparse nature of outcome rewards in RLVR settings, where only a single reward signal is provided after the entire response is generated. This sparsity severely hinders credit assignment and gradient propagation in Reinforcement Learning (RL)-based policy learning, which relies on attributing feedback to individual tokens or actions within the output. Value-model-based methods address this challenge by learning explicit value models, but at the cost of increased model complexity and computational overhead. In contrast, value-model-free methods avoid this complexity by forgoing explicit value prediction, but suffer from high variance in their Monte Carlo estimates, often leading to issues such as advantage collapse, training instability, and degraded performance.
@@
This challenge, rooted in the inherently sparse nature of RLVR feedback, highlights fundamental limitations of the existing RL-based paradigm. It motivates the development of alternative policy optimization strategies that can leverage direct supervision to enable more stable and efficient learning in RLVR settings. To this end, we propose a novel framework that recasts the RLVR problem as a supervised learning task (see Figure 1). Instead of relying on sparse outcome rewards to optimize the policy via RL, we directly interpret the outcome reward as a supervision signal and train a score function, parameterized by the policy itself, to predict this reward via a cross-entropy loss.
Through a detailed gradient analysis, we show that this supervised learning formulation not only recovers the standard policy gradient update but also implicitly couples the actor (policy improvement) and critic (reward estimation) components within a single update. Unlike conventional actor critic methods that maintain a separate value function estimator, our approach adopts a shared parameterization, allowing both roles to be updated jointly via a unified learning signal. This tight coupling eliminates the temporal mismatch between reward estimation and policy updates, thereby improving training efficiency. Moreover, in contrast to value-model-free methods, which rely solely on high-variance Monte Carlo estimates of relative advantages for policy updates, our method leverages prediction-error-based learning signals, providing more stable and reliable guidance for optimization. We refer to our method as PACS (imPlicit Actor Critic coupling via a Supervised learning framework). It offers a principled and efficient training paradigm that unifies policy learning and reward estimation within a coherent supervised learning framework.
To assess the effectiveness of our method, we conduct experiments on four challenging mathematical reasoning tasks and compare it against the state-of-the-art RLVR methods, including PPO and GRPO. The results demonstrate that PACS significantly outperforms baselines. For instance, PACS achieves a pass@256 rate of 59.78% on AIME 2025, representing improvements of 13.32 and 14.36 points over PPO and GRPO. Ablation studies confirm the critical importance of the weighting mechanism, while training dynamics analysis reveals that PACS achieves superior exploration-exploitation balance.
@@
  • We propose PACS, a novel RLVR approach that reframes the problem as supervised learning: motivated by the sparsity of outcome rewards, we treat them as target labels and train the policy via cross-entropy loss.
  • We demonstrate that the proposed loss inherently captures policy gradient updates and implicitly unifies actor (policy improvement) and critic (reward estimation) through shared parameterization, enabling efficient and stable policy optimization.
  • Extensive experiments on challenging mathematical reasoning benchmarks show that PACS outperforms state-of-the-art RLVR methods, achieving superior reasoning performance, maintaining healthy policy entropy for exploration, and exhibiting improved training efficiency.
@@ 2 Related Work
Reasoning Models.
The emergence of reasoning models, referring to LLMs explicitly trained or optimized for multi-step chain-of-thought inference, has been catalyzed by the release of OpenAI-o1 (Jaech et al., 2024), which achieves strong performance on challenging reasoning tasks. DeepSeek-R1 (Guo et al., 2025) further highlights the power of verifiable reward signals by using GRPO (Shao et al., 2024) with rule-based outcome rewards to guide effective policy optimization. Building on this foundation, QWQ (Qwen-Team, 2025) introduces a two-stage RL framework: a “cold-start” phase to bootstrap reasoning ability, followed by a brief phase of general RL finetuning to expand the model’s problem-solving proficiency. Most recently, Gemini 2.5 Pro (Comanici et al., 2025) integrates human feedback, retrieval-augmented generation, and multimodal perception, enabling reasoning grounded in diverse information sources beyond text. Together, these advances illustrate the evolution of RL-enhanced reasoning models from single-domain systems to general-purpose, multimodal agents.
Reinforcement Learning with Verifiable Reward (RLVR).
RLVR has become a widely adopted strategy for enhancing the reasoning capabilities of LLMs in domains with clearly verifiable correctness, such as mathematics and programming  (Guo et al., 2025; Team et al., 2025; Ma et al., 2025; Chu et al., 2025). Recent work has explored a wide range of RLVR techniques, which can be broadly categorized as value-model-based or value-model-free. Value-model-based methods, such as PPO (Schulman et al., 2017), VinePPO (Kazemnejad et al., 2024), and VAPO (Yue et al., 2025), explicitly learn a value function to estimate the expected cumulative reward, providing stable training signals at the cost of additional computational overhead. On the other hand, value-model-free methods, such as GRPO (Shao et al., 2024), REINFORCE++ (Hu et al., 2025), and DAPO (Yu et al., 2025), avoid explicit value modeling by relying on Monte Carlo-based advantage estimation. While these approaches reduce modeling complexity, they often suffer from high gradient variance, which undermines training stability and overall performance.
@@ 3 Methodology
In this section, we first establish the preliminaries for RLVR, and then introduce PACS, a novel RLVR framework that achieves imPlicit Actor–Critic coupling through a Supervised learning framework. We present the overall framework, followed by a detailed gradient analysis that illustrates the effectiveness of our approach, and conclude with practical design considerations to instantiate the implementation.
@@ 3.1 Preliminaries
3.2 Recasting RLVR within a Supervised Learning Framework
In the RLVR framework, LLMs generate responses to input queries, and verifiable outcome rewards that indicate the correctness of the final outcomes are provided at the end of responses. A key characteristic of this setting is that the reward is delivered only once, after the entire response is generated, reflecting the overall quality of the output. Inspired by this observation, we propose an alternative to conventional RL-based optimization: rather than policy learning through RL from sparse reward signals, we recast the problem as a supervised learning task in which the model is trained to directly predict the outcome reward.
@@ 3.2 Recasting RLVR within a Supervised Learning Framework
Concretely, instead of using the reward $R(q,o)\in\{0,1\}$ associated with a query-output pair $(q,o)$ to guide policy optimization via RL, we consider $(q,o)$ as input data and treat $R(q,o)$ as the corresponding label. The objective is to learn a mapping $f(q,o)\rightarrow R$ that accurately predicts the reward. Specifically, this problem can be cast as a classification task where rewards serve as binary labels (e.g., correct vs. incorrect). To this end, we define the learning objective optimized via a standard binary cross-entropy loss as follows:
@@ 3.3 Gradient Analysis of Objective Function
We now derive the inner gradient term $\nabla_{\theta}l(q,o;\pi_{\theta})$. Applying the chain rule yields:
$\nabla_{\theta}l(q,o;\pi_{\theta})=\left[R(q,o)\frac{\sigma^{\prime}\big{(}\psi(q,o;\pi_{\theta})\big{)}}{\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}}+\big{(}1-R(q,o)\big{)}\frac{-\sigma^{\prime}\big{(}\psi(q,o;\pi_{\theta})\big{)}}{1-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}}\right]\nabla_{\theta}\psi(q,o;\pi_{\theta})$ (4)
Recalling that the derivative of the sigmoid function is $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, we simplify the gradient term as follows:
$\nabla_{\theta}l(q,o;\pi_{\theta})=\Big{[}R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{]}\nabla_{\theta}\psi(q,o;\pi_{\theta})$ (5)
Substituting Equation 5 into the expression for $\nabla_{\theta}\mathcal{L}(\theta)$, we obtain the final form of the gradient:
$\nabla_{\theta}\mathcal{L}(\theta)=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\bigg{[}\underbrace{\big{[}l(q,o;\pi_{\theta})\big{]}\nabla_{\theta}\log\pi_{\theta}(o|q)}_{\text{{{Actor}}: policy improvement}}+\underbrace{\Big{(}R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}\nabla_{\theta}\psi(q,o;\pi_{\theta})}_{\text{{{Critic}}: reward estimation}}\bigg{]}$ (6)
In the above gradient expression, the first term corresponds to a standard policy gradient update, weighted by the per-sample cross-entropy loss. The second term serves as a reward estimation correction, adjusting the score function $\psi$ to better align the predicted reward $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ with the ground-truth outcome. Notably, this formulation unifies the Actor and Critic updates within a single gradient step and shared parameter space, enabling tighter coupling and more efficient learning.
@@
As shown in Equation 6, our formulation enables the learned model $\pi_{\theta}$ to simultaneously fulfill two roles:
@@
  • Critic: It estimates the quality of each sampled output using the score function $\psi(q,o;\pi_{\theta})$, where the sigmoid-transformed value $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ represents the predicted probability of correctness and serves as the predicted reward.
In the Actor update, the per-sample cross-entropy loss term $l(q,o;\pi_{\theta})$ quantifies the alignment between the predicted reward $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ and the ground-truth reward $R(q,o)$. This term serves as a dynamic weight in the policy gradient, modulating the magnitude of updates based on prediction accuracy. In the Critic update, the residual term $R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ captures the discrepancy between the predicted and actual reward. This prediction error serves as a direct supervised learning signal, driving the score function $\psi(q,o;\pi_{\theta})$ to better predict the outcome reward. Notably, improvements in the Critic directly influence future Actor updates by refining the weighting term $l(q,o;\pi_{\theta})$.
In general, the gradient decomposition in Equation 6 splits the gradient into two components: the policy gradient update (Actor) and the supervised reward estimation updates (Critic). Crucially, both components operate within a single policy model $\pi_{\theta}$, enabling implicit coupling between Actor and Critic without the need for separate networks or alternating update schedules. This unified design facilitates efficient training while leveraging the strengths of both policy gradient and supervised learning paradigms.
@@ 3.4 Score Function Instantiation
To instantiate the score function $\psi(q,o;\pi_{\theta})$, we require a mechanism to quantify how good a sampled response $o$ is for a given query $q$. In our work, we interpret and implement $\psi(q,o;\pi_{\theta})$ as an advantage-like function, which measures the relative quality of a response conditioned on the query within a group. Unlike value-model-free methods that estimate advantages using ground-truth rewards, we define $\psi(q,o;\pi_{\theta})$ directly in terms of the policy model $\pi_{\theta}$, enabling direct policy optimization. To compute this advantage efficiently and with minimal implementation overhead, we adopt the REINFORCE Leave-One-Out (RLOO) estimator (Kool et al., 2019; Ahmadian et al., 2024). RLOO provides an unbiased estimate of the relative advantage by comparing each sampled output to others generated for the same query within a batch. Specifically, for each query $q$, we sample a set of $G$ candidate responses $o_{1},o_{2},\dots,o_{k}$}. For each output $o_{i}$, the RLOO-based advantage-like score function is defined as:
@@
where $\hat{r}(q,o_{i};\pi_{\theta})$ is a reward proxy based on log-probability ratios, following prior work (Rafailov et al., 2023):
@@
This formulation preserves the supervised learning structure of our original objective while incorporating relative comparisons between sampled outputs, yielding richer and more informative learning signals [^1] A rule-based reward function is employed, assigning a reward of $1$ to correct answers and 0 to incorrect ones.. The overall PACS framework is illustrated in Figure 2.
3.4.1 Loss Function Analysis.
@@
To gain deeper insight into the learning dynamics, we analyze the loss function in Equation 9. When a sampled output $o_{i}$ is correct (i.e., $R(q,o_{i})=1$), the loss reduces to the term $\log(\sigma(\psi(q,o_{i};\pi_{\theta})))$. Minimizing the loss in this case corresponds to maximizing $\psi(q,o_{i};\pi_{\theta})$, which in turn requires increasing the estimated reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$. Since $\hat{r}$ is defined as a log-probability ratio, this implies increasing $\pi_{\theta}(o_{i}|q)$, i.e., assigning higher probability to generating the correct response. In this way, the model learns to favor correct outputs. Conversely, when a sampled output $o_{i}$ is incorrect (i.e., $R(q,o_{i})=0$), the loss reduces to the term $\log(1-\sigma(\psi(q,o_{i};\pi_{\theta})))$. Minimizing the loss requires decreasing $\psi(q,o_{i};\pi_{\theta})$, which again corresponds to lowering the reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$. This is achieved by reducing $\pi_{\theta}(o_{i}|q)$, discouraging the generation of incorrect responses. Thus, the model learns not only to generate correct outputs, but also to avoid producing incorrect ones.
3.4.2 Handling Data Imbalance.

### v2

@@ Abstract
Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) have empowered large language models (LLMs) to tackle challenging reasoning tasks such as mathematics and programming. Despite its promise, the RLVR paradigm poses significant challenges, as existing methods often suffer from sparse reward signals and unstable policy gradient updates, inherent to RL-based approaches. To address the challenges, we propose PACS, a novel RLVR framework that achieves imPlicit Actor Critic coupling via a Supervised learning framework. By treating the outcome reward as a predictable label, we reformulate the RLVR problem into a supervised learning task over a score function parameterized by the policy model and optimized using cross-entropy loss. A detailed gradient analysis shows that this supervised formulation inherently recovers the classical policy gradient update while providing more stable and efficient training. Extensive experiments demonstrate that PACS significantly outperforms strong open-source models and RLVR baselines, yielding substantial average gains of +8.26% (4B) and +9.57% (8B) over base models offering a promising avenue for LLMs post-training with verifiable rewards. Our code and data are available as open source at https://github.com/ritzz-ai/PACS.
@@ 1 Introduction
Recent advancements like OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025) have revolutionized complex reasoning by scaling test-time compute to enable longer Chains of Thought (CoT). This progress significantly improves performance in mathematics (Shao et al., 2024; Liu et al., 2025a; Zuo et al., 2025) and programming (Lambert et al., 2024; Cui et al., 2025; Zhao et al., 2025). These achievements are largely driven by Reinforcement Learning with Verifiable Rewards (RLVR), which empowers LLMs to self-improve using verifiable outcome rewards.
Existing RLVR methods generally fall into two categories: value-model-based (e.g., PPO (Schulman et al., 2017), VAPO (Yue et al., 2025)) and value-model-free (e.g., GRPO (Shao et al., 2024; Guo et al., 2025), DAPO (Yu et al., 2025)). While both classes have demonstrated remarkable success, they face fundamental challenges stemming from the sparse nature of outcome rewards in RLVR settings, where only a single reward signal is provided after the entire response is generated.. Value-model-based approaches mitigate sparsity via explicit value modeling but incur high computational overhead. Conversely, value-model-free methods leverage Monte Carlo estimation to avoid this cost but suffer from high variance, often leading to advantage collapse and training instability. This challenge, rooted in the inherently sparse nature of RLVR feedback, highlights fundamental limitations of the existing RL-based paradigm. It motivates the development of alternative policy optimization strategies that can leverage direct supervision to enable more stable and efficient learning in RLVR settings.
To address the limitations of the RL-based paradigm rooted in sparse feedback, we propose PACS (imPlicit Actor Critic coupling via a Supervised learning framework). We recast RLVR as a supervised learning task (as illustrated in fig. 1) by treating outcome rewards as labels to train a policy-parameterized score function via cross-entropy loss. Gradient analysis reveals that this formulation inherently recovers the standard policy gradient while implicitly coupling the actor and critic through shared parameterization. This design eliminates the temporal mismatch introduced by separate value estimators and mitigates the high-variance issues associated with Monte Carlo estimates by providing stable prediction-error signals for optimization. Moreover, it reduces compute overhead compared to value-model-based training, and attenuates extreme updates that cause entropy collapse via bounded supervised residual. Overall, PACS offers a principled and efficient training paradigm that unifies policy learning and reward estimation within a coherent supervised learning framework.
@@
Experimental results on five benchmarks demonstrate PACS’s superiority. It outperforms strong open-source models like LIMO and Eurus-2-7B with a 63.32% $\text{pass@}1$ accuracy. Notably, PACS 4B model (45.63% on AIME 2025) surpasses the significantly larger Qwen3-14B (36.22%). Against RLVR baselines, PACS establishes a decisive lead, particularly on complex tasks: on BeyondAIME, PACS-8B achieves 64.02% ($\text{pass@}64$), exceeding GRPO (57.64%) and PPO (53.80%) by approximately 6.4 and 10.2 points, respectively. Furthermore, PACS mitigates entropy collapse, achieving significantly higher solution diversity than GRPO (0.0392 vs. 0.0249).
@@
  • We propose PACS, a novel RLVR framework that recasts sparse outcome rewards as supervised signals, effectively training the policy via cross-entropy loss.
  • We theoretically demonstrate that the proposed loss inherently captures policy gradient updates and implicitly unifies actor and critic through shared parameterization, enabling efficient and stable policy optimization.
  • Extensive experiments on challenging mathematical reasoning benchmarks show that PACS consistently outperforms state-of-the-art RLVR methods, achieving superior reasoning performance.
  • Generalization and diversity analyses demonstrate that PACS possesses strong transfer capabilities on out-of-domain tasks and effectively mitigates entropy collapse, fostering diverse reasoning paths that are instrumental to the model’s superior performance.
@@ 2 Related Work
Reasoning Models. The emergence of reasoning models was catalyzed by OpenAI-o1 (Jaech et al., 2024), which demonstrated that reasoning performance can scale significantly via test-time compute. This paradigm was further advanced by DeepSeek-R1 (Guo et al., 2025), which utilized GRPO (Shao et al., 2024) to show that rule-based rewards can elicit self-correction and reflection without dense human supervision. This paradigm has been extended through the Qwen family (Team, 2025b; Qwen et al., 2025; Yang et al., 2025) and the Kimi series (Team et al., 2025c; b), which have brought high-performance reasoning to the ecosystem. Most recently, Gemini family (Comanici et al., 2025) and the Llama-Nemotron series (Grattafiori et al., 2024; Bercovich et al., 2025) have integrated multimodal perception and instruction-following into the reasoning loop, establishing a new state-of-the-art for general-purpose agents.
Reinforcement Learning with Verifiable Reward (RLVR). RLVR has become a widely adopted strategy for enhancing the reasoning capabilities of LLMs in domains with clearly verifiable correctness, such as mathematics and programming  (Guo et al., 2025; Ma et al., 2025; Chu et al., 2025; Yan et al., 2025; Zheng et al., 2025; Hao et al., 2025). Recent work has explored a wide range of RLVR techniques, which can be broadly categorized as value-model-based or value-model-free. Value-model-based methods, such as PPO (Schulman et al., 2017), VinePPO (Kazemnejad et al., 2024), and VAPO (Yue et al., 2025), explicitly learn a value function to estimate the expected cumulative reward, providing stable training signals at the cost of additional computational overhead. On the other hand, value-model-free methods, such as GRPO (Shao et al., 2024), REINFORCE++ (Hu et al., 2025), and DAPO (Yu et al., 2025), avoid explicit value modeling by relying on Monte Carlo advantage estimation. While these approaches reduce modeling complexity, they often suffer from high gradient variance, which undermines training stability and performance.
@@ 3 Methodology
In this section, we first establish the preliminaries for RLVR before introducing PACS, a novel framework achieving imPlicit Actor–Critic coupling via Supervised learning. We then detail the overall framework, validate its effectiveness through gradient analysis, and conclude with practical implementation designs.
@@ 3.1 Preliminaries
3.2 Recasting RLVR via Supervised Learning
In the RLVR framework, LLMs receive verifiable outcome rewards only after generating a complete response. Inspired by this observation, we propose an alternative to conventional RL-based optimization: rather than policy learning through RL from sparse reward signals, we recast the problem as a supervised learning task in which the model is trained to directly predict the outcome reward.
Concretely, instead of using the reward $R(q,o)\in\{0,1\}$ associated with a query-output pair $(q,o)$ to guide policy optimization via RL, we consider $(q,o)$ as input data and treat $R(q,o)$ as the corresponding label. The objective is to learn a mapping $f(q,o)\rightarrow R$ that accurately predicts the reward. Specifically, this problem can be cast as a classification task where rewards serve as binary labels (e.g., correct vs. incorrect). To this end, we define the learning objective via a standard binary cross-entropy loss as follows:
@@ 3.3 Gradient Analysis of Objective Function
We now derive the inner gradient term $\nabla_{\theta}\ell(q,o;\pi_{\theta})$. Applying the chain rule and recalling that the derivative of the sigmoid function is $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, we simplify the gradient term as follows:
$\nabla_{\theta}\ell(q,o;\pi_{\theta})=\Big[R(q,o)-\sigma\big(\psi(q,o;\pi_{\theta})\big)\Big]\nabla_{\theta}\psi(q,o;\pi_{\theta}).$ (4)
Substituting eq. 4 into the expression for $\nabla_{\theta}\mathcal{L}(\theta)$, we obtain the final form of the gradient:
$\begin{aligned} \nabla_{\theta}\mathcal{L}(\theta)=&-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\\ &\bigg[\underbrace{\big[\ell(q,o;\pi_{\theta})\big]\nabla_{\theta}\log\pi_{\theta}(o|q)}_{\text{{{Actor}}: policy improvement}}+\underbrace{\Big(R(q,o)-\sigma\big(\psi(q,o;\pi_{\theta})\big)\Big)\nabla_{\theta}\psi(q,o;\pi_{\theta})}_{\text{{{Critic}}: reward estimation}}\bigg].\end{aligned}$ (5)
In the above gradient expression, the first term corresponds to a standard policy gradient update, weighted by the per-sample cross-entropy loss. The second term serves as a reward estimation correction, adjusting the score function $\psi$ to align the predicted reward $\sigma\big(\psi(q,o;\pi_{\theta})\big)$ with the ground-truth outcome. Notably, this formulation unifies the Actor and Critic updates within a single gradient step and shared parameter space.
@@
As shown in eq. 5, our formulation enables the learned model $\pi_{\theta}$ to simultaneously fulfill two roles:
@@
  • Critic: It estimates output quality using the score function $\psi(q,o;\pi_{\theta})$, where the sigmoid-transformed value $\sigma\big(\psi(q,o;\pi_{\theta})\big)$ represents the predicted probability of correctness and serves as the predicted reward.
The Actor update is modulated by the per-sample cross-entropy loss $\ell(q,o;\pi_{\theta})$, which weights the policy gradient according to the alignment between predicted and ground-truth rewards.
The Critic update is driven by the residual $R(q,o)-\sigma\big(\psi(q,o;\pi_{\theta})\big)$, providing a supervised learning signal that improves reward prediction. Improving the Actor changes the score $\psi$ and thus yields a tighter supervised target for the Critic; in turn, the Critic improvements feed back to directly affect future Actor updates through the shared weighting term. This unified design enables implicit Actor Critic coupling without separate networks or alternating update schedules and facilitates efficient training while leveraging the strengths of both policy gradient and supervised learning paradigms.
@@ 3.4 Score Function Instantiation
To instantiate the score function $\psi(q,o;\pi_{\theta})$, we require a mechanism to quantify the quality of a sampled response $o$ for a given query $q$. We interpret $\psi(q,o;\pi_{\theta})$ as an advantage-like function that measures the relative quality of a response within a group of samples.
Such a formulation is particularly appropriate for our setting because it centers the score distribution with respect to the set of samples while preserving its full range. As a result, it is well suited with a sigmoid mapping, yielding outputs confined to the $[0,1]$ interval.
Unlike value-model-free methods that estimate advantages using ground-truth rewards, we define $\psi(q,o;\pi_{\theta})$ in terms of the policy model $\pi_{\theta}$, enabling direct policy optimization. To compute this advantage efficiently, we adopt the REINFORCE Leave-One-Out (RLOO) estimator (Kool et al., 2019; Ahmadian et al., 2024). RLOO provides an unbiased estimate of the relative advantage by comparing each sampled output to others generated for the same query. For each query $q$, given a set of $G$ candidate responses $\{o_{1},o_{2},\dots,o_{k}\}$, the RLOO-based advantage-like score function for each output $o_{i}$ is defined as:
@@
following prior work (Rafailov et al., 2023; Cui et al., 2025), where $\hat{r}(q,o_{i};\pi_{\theta})$ is a reward proxy based on log-probability ratios:
@@
Loss Function Analysis.
@@
For a correct output $o_{i}$ (i.e., $R(q,o_{i})=1$), the loss reduces to $\log(\sigma(\psi(q,o_{i};\pi_{\theta})))$. Minimizing this term maximizes $\psi(q,o_{i};\pi_{\theta})$, thereby increases the reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$ and, equivalently, the probability $\pi_{\theta}(o_{i}|q)$ of generating the correct output. Conversely, for an incorrect output $o_{i}$ (i.e., $R(q,o_{i})=0$), the loss becomes $\log(1-\sigma(\psi(q,o_{i};\pi_{\theta})))$. Minimizing it decreases $\psi(q,o_{i};\pi_{\theta})$, which in turn lowers $\hat{r}(q,o_{i};\pi_{\theta})$ and reduces the probability $\pi_{\theta}(o_{i}|q)$ of generating the incorrect output. Thus, the model learns not only to generate correct outputs, but also to avoid producing incorrect ones.
Model MATH 500 AMC 23 AIME 2024 AIME 2025 BeyondAIME Avg.
Closed-source Models
Claude Opus 4.1 - 87.5 - 78.0 - -
Gemini 2.5 pro 96.7 100.0 88.7 86.7 58.8 86.18
Open-Source Model
Qwen3-4B 91.01 82.30 46.48 34.71 17.33 54.37
Qwen3-8B 89.94 81.74 46.98 33.83 16.28 53.75
Qwen3-14B 92.07 84.88 49.06 36.22 18.59 56.16
Deepthought-8B 45.07 16.58 1.82 0.44 0.76 12.93
Eurus-2-7B-PRIME 82.07 63.65 18.39 13.88 6.67 36.93
LIMO 91.00 80.61 40.86 31.82 15.98 52.05
Marco-o1 70.48 45.12 9.22 6.95 2.77 26.91
OpenThinker3-7B 88.90 72.44 36.95 26.98 12.33 47.52
Sky-T1-7B 85.38 69.59 20.86 20.47 8.54 40.97
PACS-4B 94.80 90.45 55.10 45.63 27.16 62.63
$\Delta(\text{Qwen3-4B})$ $\uparrow$3.79 $\uparrow$8.15 $\uparrow$8.62 $\uparrow$10.92 $\uparrow$9.83 $\uparrow$8.26
PACS-8B 95.09 88.69 57.58 46.38 28.86 63.32
$\Delta(\text{Qwen3-8B})$ $\uparrow$5.15 $\uparrow$6.95 $\uparrow$10.6 $\uparrow$12.55 $\uparrow$12.58 $\uparrow$9.57
Table 1: Comparison of $\text{pass@}1$ accuracy (%) on mathematical reasoning benchmarks. “Avg.” indicates the average performance across all tasks. Bold numbers indicate the best performance. Underlined numbers indicate the second best. The values in green represent the performance gain ($\Delta$) over the model.
Overall, this formulation preserves the supervised learning structure of our original objective while incorporating relative comparisons between sampled outputs, yielding more informative learning signals [^1] A rule-based reward function is employed, assigning a reward of $1$ to correct answers and 0 to incorrect ones.. The overall PACS framework is illustrated in fig. 2.
Mitigating Entropy Collapse.
With the score function instantiated in a policy-dependent form, the gradient of the loss can be expressed as an advantage-weighted policy gradient (Appendix A.1). Unlike conventional RL, where positive rewards can push the policy toward overly deterministic behavior, our gradient (eq. 5) is modulated by the residual $R(q,o)-\sigma(\psi(q,o;\pi_{\theta}))$. As predictions approach the ground truth, the residual vanishes, naturally attenuating updates and preventing over-concentration of probability mass, thereby preserving a diverse set of responses. We provide a formal proof in Appendix A.2 showing that this mechanism implicitly imposes a structural constraint that regularizes the policy towards a Gibbs distribution, fundamentally preventing mode collapse.
3.5 Practical Considerations: Handling Data Imbalance